### v3

@@ 4.4.2 Step-size renewal mechanism
5 Experiments
In this section, we first compare our methods with the existing works using ResNet on CIFAR [27] and ImageNet-1k [9], then conduct ablation studies to verify different aspects of our methods, including (1) each proposed technology’s verification, (2) our methods’ extension to spiking self-attention [56], (3) mitigating the "temporal constraint" of the prior arts [39] on the event dataset CIFAR10-DVS [28] and DVS-GESTURE [2], (4) extending to the spiking audio task SHD [8], (5) the learned bit allocation’s visualization, and (6) comparing the mixed-precision ANNs. All implementation details are attached in the appendix.
5.1 Comparisons with existing works
5.1.1 CIFAR
We apply our methods to ResNet20 [18] and ResNet18 [29] to compete with the advanced SNNs. As shown in Table 1, we first compare our methods with those using extra learning- or module-based model optimization [29, 11, 55], and achieve superior accuracy using much lower memory (bit budget) and computation (S-ACE). In the case of ResNet-18 with target W/S/T being 1/1/1, our methods can also compress the model sizes to one-tenth of their original dimensions, while achieving at most 2% and 3.74% higher model accuracy on CIFAR-10 and CIFAR-100, respectively. Such a reduction of storage requirements would make SNNs more compatible with on-chip computing architectures [1, 34]. Comparing with Ternary Spike ResNet20 [18] that adopts 2-bit bidirectional spike, we can outperform it by at least 1.19% and 5.88% accuracy increase on CIFAR-10 and -100, respectively, while using average 1.32- and 1.58-bit spike and reducing 91.37% and 90.59% model sizes. Then, we compare with Multi-bit ResNet20 [48], whose memory and computation respectively increase to huge 256 bit budgets and 660.48 G S-ACEs. While, our model can achieve not only 1.22% and 3.39% accuracy increase on CIFAR-10 and CIFAR-100 but also two magnitudes of memory and computation savings. Finally, compared with plain uniform quantization (U-quant.), our methods can maximize the model accuracy while minimizing memory and computation requirements. The advantages of our approaches in terms of memory usage, computational efficiency, and storage requirements are consistently demonstrated. Additionally, since we adopt temporal squeezing as shown in Figure 2 and Equation 13, the temporal dimension can be compressed to 1 time-step before operating convolution. That is why S-ACE can further decrease when "W/S/T=222" and more elaborations are in the appendix for better understanding. In conclusion, by comparing with state-of-the-art methods on CIFAR-10 and CIFAR-100, we demonstrate significant improvements in memory efficiency, computational requirements, and storage consumption.
5.1.2 ImageNet
We further apply our methods to SEW-ResNet34 to demonstrate superiority on the up-scaled dataset. As listed in Table 2, compared with those using improved learning techniques or special modules [36, 55, 29, 22], we can achieve magnitudes of memory and computation savings without using any of their tricks. Attention SNN [52] has comparable S-ACE (56.73 G) with ours (30.74 G and 54.69 G), but its memory efficiency, model accuracy, and storage requirement lag far behind those of our methods. We can achieve at least 1.42% increased accuracy while reducing 45.81% S-ACE and 83.86% model size. Besides, our models would not require sophisticated attention modules [52] as they do, which indicates a more straightforward inference dataflow and a better computing flow mapping from software to hardware. Furthermore, we can save 91.2% memory (5.61 v.s. 64) and 85.83% computation (30.74 G v.s. 217 G) while achieving a higher accuracy (70.57% v.s. 70.13%) when comparing with the quantized SNN [39]. As the bit budget increases, the model accuracy can be advanced to 72.82% with low memory and computation overheads, outperforming full-precision SNNs [18, 15, 11]. Conclusively, scaling our model to the large ImageNet-1k dataset further demonstrates the effectiveness of the proposed methods in improving the efficiency and accuracy of SNNs.
@@ 5.1.2 ImageNet
5.2 Ablation study
We first continue to use ResNet20 and CIFAR10 to conduct ablation studies on the proposed techniques. Then, we migrate the proposed method to a different architecture, Spikformer [56], to further verify effectiveness. On the extension to the event dataset, we leverage our methods to mitigate the "temporal constraint" issue, which troubles the prior art [39]. We also extend our methods to the spiking audio task SHD [8] by incorporating the axonal delay mechanism [40, 20] Furthermore, we provide visualizations of the learned bit allocation as empirical evidence and statistically analyze the results. Finally, we compare our bit-adaptive SNNs with the ANN baselines to show the superiority of SNNs in scenarios of ultra-low-energy computation. Notably, since our work targets building efficient and accurate SNNs, we mainly focus on the effectiveness of the proposed approaches under the low bit budget settings and conduct ablation studies in this background.
@@ 5.2.1 Effectiveness of adaptive bit width and the renewal mechanism
We first perform the plain uniform quantization on SNNs (U-quant.) as the baselines, following [39]. As shown in Table 3, the proposed adaptive bit-width technology consistently outperforms U-quant., managing to allocate valid bit widths to different layers and improving the model accuracy to a higher level. Furthermore, we add the proposed renewal mechanism only to the spike activation ($V_{th,l}^{1}$, Act.-only), or only to the weight ($S_{q}^{l}$, Weight-only), or to the bilateral case ($V_{th,l}^{1}$ and $S_{q}^{l}$). The results coherently show the renewal mechanism can further improve the outcomes of the bit-adaptive models, mitigating the step-size mismatch issue. As described before, SNN’s activation would accumulate quantization errors along the temporal dimension. That is why Act.-only renewal is effective enough to produce the best results. Bilateral renewal seconds due to a slightly lower bit budget. For simplicity, we adopt the Act.-only renewal as the main method by default.
In addition, another interesting finding is that the results of the target W/S/T=4/4/1 are better than those of the target W/S/T=4/4/2. This may imply that more temporal information contribute limited improvement or even cause over-fitting to the processing of static data when spatial information is efficiently provided and processed.
@@
Target U-quant. Bit-adaptive Bit-adaptive + renewal Origin [56] W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 1/1/1 79.53 1.44/1.48/1.0 95.37 1.53/1.5/1.0 95.47 1/2/1 89.3 1.47/1.99/1.0 95.68 1.47/2.01/1.0 95.54 2/2/1 89.68 1.88/2.03/1.0 95.80 2.02/2.01/1.0 95.84 2/3/1 94.11 2.00/2.94/1.0 95.93 2.15/2.95/1.0 96.04 3/3/1 94.21 2.93/3.01/1.0 95.87 2.89/2.99/1.0 95.99 16/1/4 95.19 U-quant. denotes plain uniform quantization, following [39].
@@ 5.2.2 Importance of the refined spiking neuron
We further verify the effectiveness of the refined spiking neuron. Key improvements over prior multi-bit spiking neurons [18, 48] include: (1) refining Equation 5 to Equation 15 by replacing flooring with rounding, and (2) adopting the IF neuron ($\tau=1$) instead of the default LIF neuron ($\tau=2$) [14]. In Table 4, the results consistently show that the original multi-bit neuron would impair the quantized model’s robustness, whereas our refinement method and renewal mechanism steadily improve the model performance to better states. This enables the development of high-performance and low-overhead SNNs.
Another insight might also be drawn. In the field of deep learning, the value quantizer regulates information widths of features and consequently reduce the redundant data movement and computing operations in the following layer. The firing mechanism in the LIF model emulates biological neurotransmission by transforming chemical stimuli into electrical impulses, achieving energy-efficient information transfer [17]. Therefore, the firing mechanism is the key to generating the efficient value representation, i.e., low-bit value, to achieve energy-efficient computing as the value quantizer does. Previous works [46, 23, 19] on direct-trained SNNs directly adopted the mathematical formulation of the LIF model without adapting it to the intrinsic characteristics of deep learning computations, which potentially compromises the network performance. After introducing the concept of quantization error and adhering to the principle of minimizing quantization errors, we refined the LIF model, thereby achieving further performance enhancements in SNNs. In conclusion, although SNN is a bio-inspired model, reducing the quantization error is still important in maintaining SNN’s good model performance.
@@ 5.2.3 Validation on spiking self-attention architectures
Migrating our methods to Spikformer is non-trivial because the self-attention operation is a different computing paradigm where a new quantizable tensor attenatten (see Figure 2(a)) is produced. We would like to leave room for future work. Here, we naively apply our methods to Spikformer as depicted in Figure 2 only to validate the migratability and effecitveness of the proposed methods, and avoid competing with those state-of-the-art works that are based on sophisticated spiking self-attention architectures. We adopt the default Spikformer-4-384 [56] on CIFAR10. In Table 5, our bit-adaptive model surpasses the original Spikformer, while plain uniform quantization fails. Besides, adding the renewal mechanism further lifts accuracy.
@@ 5.2.4 Mitigating the temporal constraint
We further test the migratability of the proposed methods on the event datasets. We use the same default Spikformer and settings as [39] without data augmentation. As shown in Table 6, "temporal constraint" refers to the dramatic accuracy decrease of event datasets as temporal length declines to the extremely low time-step [39]. However, in our work, such a phenomenon disappears and accuracy is well maintained even in the case of $T=1$. This implies that our methods can better extract spatial information even in sparse event data. Meanwhile, we are in line with prior findings that temporal length would maintain a notable effect on accuracy for event datasets. Moreover, compared with the original Spkiformer, our methods can achieve higher accuracy with lower bit budgets and time-steps, proving the effectiveness on event data.
Futhermore, another mainstream event-based dataset DVS-GESTURE [2] is smaller and easier for models to converge, and the "temporal constraint" issue does not manifest. In this case, our model can still significantly strengthen the temporal processing ability of SNNs. As listed in Table 7, our model can achieve consistently better accuracy while using shorter time lengths and fewer bit budgets when compared with the baseline models. In the case of W/S/T = 1.17/2.18/6.0, our model can achieve advanced 99.31 % top-1 accuracy. To the best of our knowledge, this should be a state-of-the-art result of SNNs on the dynamic dataset.
@@
Origin [39] Quantized Spkiformer [39] Our Spkiformer W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 16/1/16 80.7 1/1/16 79.8 1.11/2.00/6.0 80.9 1/2/8 79.3 1.28/2.00/4.0 80.0 1/4/4 63.1 1.33/4.00/3.0 80.1 1/8/2 43.0 1.31/4.03/2.0 78.7 1/16/1 35.8 1.31/4.03/1.0 77.4
@@
Origin [39, 56] Quantized Spkiformer [39] Our Spkiformer W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 16/1/16 98.3 1/1/16 96.67 1.17/2.18/6.0 99.31 1/2/8 98.48 1.29/2.08/4.0 98.61 1/4/4 97.35 1.32/4.00/3.0 97.92 1/8/2 96.59 1.31/4.00/2.0 97.57 1/16/1 95.45 1.37/4.00/1.0 96.18
Table 8: Extended experiments on the spiking audio dataset SHD.
Method Parameters (M) Top1 DMUC [41] 0.24 91.48 CNN [8] 1.01 92.40 RSNN [8] 1.97 83.20 Attention RSNN [51] 0.14 91.08 RadLIF [4] 3.9 94.62 DL256-SNN-Dloss [40] 0.21 93.55 DCLS [20] 0.21 95.07 Our DCLS-3.5/1/140 0.21 95.17 Our DCLS-4.4/1/140 0.21 95.54 "Ours DCLS-x/x/x" denotes the learned bit widths of W/S/T. Following the research tradition of spiking audio tasks, we only report the parameter numbers (M) instead of the model size (MB).
5.2.5 Extension to the spiking audio task
Instead of focusing solely on the event image datasets, including CIFAR10-DVS and DVS-GESTURE, we also extend our methods to the spiking audio dataset SHD [8]. However, directly applying SNNs to the spiking audio dataset will cause significant performance loss [40]. Prior arts have proposed that the axonal delay mechanism can solve this issue [40, 20]. Therefore, we utilize the dilated convolution with learnable spacings (DCLS) [20] to implement the axonal delay. Instead of directly adopting the local spike connection, which may require some modifications on the spiking neuron formulation [40], to realize axonal delays, DCLS utilizes the temporal-wise 1D convolution to implement the axonal delay mechanism. This means DCLS only operates on spiking feature maps without introducing any modification to our refined spiking neurons.
As shown in Table 8, our method not only outperforms the DCLS baseline, but also achieves the state-of-the-art (SoTA) model accuracy on the SHD dataset. This result shows our bit-adaptive SNNs are compatible with the axonal delay mechanism, and can significantly improve the SNN performance when handling the spiking audio data.
@@ 5.2.6 Statistics of bit widths
We visualize the learned bit allocation of ResNet20 with W/S/T = 1.38/1.32/1 on CIFAR10 to visually confirm the effectiveness. Notably, different from any prior arts [18, 55, 15], we add additional spiking neurons before the first layer to fully benefit the model with addition-only computation. In consistence with the previous setting, the model is initialized to W/S/T = 4/4/2. As vividly displayed in Figure 4, I) the layers with the higher proportion of features and weights tend to get lower bits thus lowering the overall memory and computation. II) The first and last layers tend to learn higher bit widths in line with their outstanding importance as reported in prior arts [7]. With I) and II) combined, we conclude our adaptive bit allocation method should be sensible and practical.
@@ 5.2.7 Additional comparisons with mixed-precision ANNs
Since we use a similar bit-width parametrization and gradient calculation to the mixed-precision ANN [42], we also need to conduct a fair comparison with mixed-precision ANNs. Here, we continue using ResNet20 on CIFAR10.
@@
Target Mixed-precision ANN [42] Our bit-adaptive SNN W/A(S)/T W/A Top-1 W/S/T Top-1 1/1/1 1.48/1.93 95.78 1.25/1.36/1.0 95.84 1/2/1 1.48/2.03 95.68 1.25/2.04/1.0 96.25 2/2/1 1.92/2.07 95.71 2.00/1.89/1.0 96.47 2/3/1 2.19/3.05 96.29 2.02/3.06/1.0 96.97 3/3/1 3.13/2.97 96.56 2.87/3.04/1.0 96.77 3/4/1 3.07/3.98 96.97 2.97/3.94/1.0 97.01 4/4/1 3.87/3.96 96.76 3.86/4.00/1.0 97.03 The results of mixed-precision ANNs [42] are produced by re-implementing their paper.
As listed in Table 9, our proposed adaptive bit allocation SNNs consistently outperform the mixed-precision ANNs under every W/S/T setting. In the case of extremely low bit widths, e.g., the target W/S/T=1/1/1, our SNN can achieve higher accuracy while using a lower bit budget. This implies that SNNs may have an advantage over ANNs in scenarios of ultra-low-energy computation, which is in line with previous studies [10, 21].
@@ 6 Discussion
Exploiting the spike perceptrons. Temporal squeezing itself is consistent with neurodynamics as it is essentially composed of a signal accumulation [17] and a division operation. A similar accumulation mechanism can also be found in the current accumulation stage of spiking neuron, such as LIF and IF. Particularly, SpikeConverter neuron [32] separates this accumulation from the spiking neuron modeling and uses a special computing pipeline to handle this accumulation. Their accumulation is identical to ours, which is proven feasible in asynchronous computing. In addition, such accumulation has also been realized and implemented by neuromorphic circuits to process the DVS signals and generate the dynamic datasets [2, 31].
@@
Using the two-stage pipeline. Although we have theoretically elaborated on the harmlessness of temporal squeezing to the asynchronous nature in the above, one may argue that the entire network’s spike flow could be interrupted by the squeezing stage, leading to longer inference latency. This concern has already been discussed and resolved by [32], where the two-stage pipeline is adopted. The first stage is spike accumulation, i.e., the temporal squeezing, the second stage is spike emission. These two separate stages can overlapped inter-layer. Therefore, using temporal squeezing only introduces very limited latency increase.

### v4

@@ 5.1.2 ImageNet
5.2 Ablation study
We first continue to use ResNet20 and CIFAR10 to conduct ablation studies on the proposed techniques. Then, we migrate the proposed method to a different architecture, Spikformer [zhou2022spikformer], to further verify effectiveness. On the extension to the event dataset, we leverage our methods to mitigate the "temporal constraint" issue, which troubles the prior art [shen2024conventional]. Furthermore, we provide visualizations of the learned bit allocation as empirical evidence and statistically analyze the results. Finally, we compare our bit-adaptive SNNs with the ANN baselines to show the superiority of SNNs in scenarios of ultra-low-energy computation. Notably, since our work targets building efficient and accurate SNNs, we mainly focus on the effectiveness of the proposed approaches under the low bit budget settings and conduct ablation studies in this background.
@@ 5.2.1 Effectiveness of adaptive bit width and the renewal mechanism
We first perform the plain uniform quantization on SNNs (U-quant.) as the baselines, following [shen2024conventional]. As shown in Table 3, the proposed adaptive bit-width technology consistently outperforms U-quant., managing to allocate valid bit widths to different layers and improving the model accuracy to a higher level. Furthermore, we add the proposed renewal mechanism only to the spike activation ($V_{th,l}^{1}$, Act.-only), or only to the weight ($S_{q}^{l}$, Weight-only), or to the bilateral case ($V_{th,l}^{1}$ and $S_{q}^{l}$). The results coherently show the renewal mechanism can further improve the outcomes of the bit-adaptive models, mitigating the step-size mismatch issue. As described before, SNN’s activation would accumulate quantization errors along the temporal dimension. That is why Act.-only renewal is effective enough to produce the best results. Bilateral renewal seconds due to a slightly lower bit budget. For simplicity, we adopt the Act.-only renewal as the main method by default.
In addition, another interesting finding is that the results of the target W/S/T=4/4/1 are better than those of the target W/S/T=4/4/2. This may imply that more repeated static information contribute limited improvement or even cause over-fitting to the processing of static data when spatial information is efficiently provided and processed.
@@
Target U-quant. Bit-adaptive Bit-adaptive + renewal Origin [zhou2022spikformer] W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 1/1/1 79.53 1.44/1.48/1.0 95.37 1.53/1.5/1.0 95.47 1/2/1 89.3 1.47/1.99/1.0 95.68 1.47/2.01/1.0 95.54 2/2/1 89.68 1.88/2.03/1.0 95.80 2.02/2.01/1.0 95.84 2/3/1 94.11 2.00/2.94/1.0 95.93 2.15/2.95/1.0 96.04 3/3/1 94.21 2.93/3.01/1.0 95.87 2.89/2.99/1.0 95.99 16/1/4 95.19 U-quant. denotes plain uniform quantization, following [shen2024conventional].
@@ 5.2.2 Importance of the refined spiking neuron
We further verify the effectiveness of the refined spiking neuron. Key improvements over prior multi-bit spiking neurons [guo2024ternary, xiao2024multi] include: (1) refining Equation 5 to Equation 15 by replacing flooring with rounding, and (2) adopting the IF neuron ($\tau=1$) instead of the default LIF neuron ($\tau=2$) [fang2023spikingjelly]. In Table 4, the results consistently show that the original multi-bit neuron would impair the quantized model’s robustness, whereas our refinement method and renewal mechanism steadily improve the model performance to better states. This enables the development of high-performance and low-overhead SNNs.
Another insight might also be drawn. In the field of deep learning, the value quantizer regulates information widths of features and consequently reduce the redundant data movement and computing operations in the following layer. The firing mechanism in the LIF model emulates biological neurotransmission by transforming chemical stimuli into electrical impulses, achieving energy-efficient information transfer [gerstner2014neuronal]. Therefore, the firing mechanism is the key to generating the efficient value representation, i.e., low-bit value, to achieve energy-efficient computing as the value quantizer does. Previous works [wu2019direct, hu2021spiking, guo2022real] on direct-trained SNNs directly adopted the mathematical formulation of the LIF model without adapting it to the intrinsic characteristics of deep learning computations, which potentially compromises the network performance. After introducing the concept of quantization error and adhering to the principle of minimizing quantization errors, we refined the LIF model, thereby achieving further performance enhancements in SNNs. In conclusion, although SNN is a bio-inspired model, reducing the quantization error is still important in maintaining SNN’s good model performance.
@@ 5.2.3 Validation on spiking self-attention architectures
Migrating our methods to Spikformer is non-trivial because the self-attention operation is a different computing paradigm where a new quantizable tensor attenatten (see Figure 2(a)) is produced. We would like to leave room for future work. Here, we naively apply our methods to Spikformer as depicted in Figure 2 only to validate the migratability and effecitveness of the proposed methods, and avoid competing with those state-of-the-art works that are based on sophisticated spiking self-attention architectures. We adopt the default Spikformer-4-384 [zhou2022spikformer] on CIFAR10. In Table 5, our bit-adaptive model surpasses the original Spikformer, while plain uniform quantization fails. Besides, adding the renewal mechanism further lifts accuracy.
@@ 5.2.4 Mitigating the temporal constraint
We further test the migratability of the proposed methods on the event datasets. We use the same default Spikformer and settings as [shen2024conventional] without data augmentation. As shown in Table 6, "temporal constraint" refers to the dramatic accuracy decrease of event datasets as temporal length declines to the extremely low time-step [shen2024conventional]. However, in our work, such a phenomenon disappears and accuracy is well maintained even in the case of $T=1$. This implies that our methods can better extract spatial information even in sparse event data. Meanwhile, we are in line with prior findings that temporal length would maintain a notable effect on accuracy for event datasets. Moreover, compared with the original Spkiformer, our methods can achieve higher accuracy with lower bit budgets and time-steps, proving the effectiveness on event data.
Futhermore, another mainstream event-based dataset DVS-GESTURE [amir2017low] is smaller and easier for models to converge, and the "temporal constraint" issue does not manifest. In this case, our model can still significantly strengthen the temporal processing ability of SNNs. As listed in Table 7, our model can achieve consistently better accuracy while using shorter time lengths and fewer bit budgets when compared with the baseline models. In the case of W/S/T = 1.17/2.18/6.0, our model can achieve advanced 99.31 % top-1 accuracy. To the best of our knowledge, this should be a state-of-the-art result of SNNs on the dynamic dataset.
@@
Origin [shen2024conventional] Quantized Spkiformer [shen2024conventional] Our Spkiformer W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 16/1/16 80.7 1/1/16 79.8 1.11/2.00/6.0 80.9 1/2/8 79.3 1.28/2.00/4.0 80.0 1/4/4 63.1 1.33/4.00/3.0 80.1 1/8/2 43.0 1.31/4.03/2.0 78.7 1/16/1 35.8 1.31/4.03/1.0 77.4
@@
Origin [shen2024conventional, zhou2022spikformer] Quantized Spkiformer [shen2024conventional] Our Spkiformer W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 16/1/16 98.3 1/1/16 96.67 1.17/2.18/6.0 99.31 1/2/8 98.48 1.29/2.08/4.0 98.61 1/4/4 97.35 1.32/4.00/3.0 97.92 1/8/2 96.59 1.31/4.00/2.0 97.57 1/16/1 95.45 1.37/4.00/1.0 96.18
@@ 5.2.5 Statistics of bit widths
We visualize the learned bit allocation of ResNet20 with W/S/T = 1.38/1.32/1 on CIFAR10 to visually confirm the effectiveness. Notably, different from any prior arts [guo2024ternary, zheng2021going, fang2021deep], we add additional spiking neurons before the first layer to fully benefit the model with addition-only computation. In consistence with the previous setting, the model is initialized to W/S/T = 4/4/2. As vividly displayed in Figure 4, I) the layers with the higher proportion of features and weights tend to get lower bits thus lowering the overall memory and computation. II) The first and last layers tend to learn higher bit widths in line with their outstanding importance as reported in prior arts [choi2018pact]. With I) and II) combined, we conclude our adaptive bit allocation method should be sensible and practical.
@@ 5.2.6 Additional comparisons with mixed-precision ANNs
Since we use a similar bit-width parametrization and gradient calculation to the mixed-precision ANN [uhlich2019mixed], we also need to conduct a fair comparison with mixed-precision ANNs. Here, we continue using ResNet20 on CIFAR10.
@@
Target Mixed-precision ANN [uhlich2019mixed] Our bit-adaptive SNN W/A(S)/T W/A Top-1 W/S/T Top-1 1/1/1 1.48/1.93 95.78 1.25/1.36/1.0 95.84 1/2/1 1.48/2.03 95.68 1.25/2.04/1.0 96.25 2/2/1 1.92/2.07 95.71 2.00/1.89/1.0 96.47 2/3/1 2.19/3.05 96.29 2.02/3.06/1.0 96.97 3/3/1 3.13/2.97 96.56 2.87/3.04/1.0 96.77 3/4/1 3.07/3.98 96.97 2.97/3.94/1.0 97.01 4/4/1 3.87/3.96 96.76 3.86/4.00/1.0 97.03 The results of mixed-precision ANNs [uhlich2019mixed] are produced by re-implementing their paper.
As listed in Table 8, our proposed adaptive bit allocation SNNs consistently outperform the mixed-precision ANNs under every W/S/T setting. In the case of extremely low bit widths, e.g., the target W/S/T=1/1/1, our SNN can achieve higher accuracy while using a lower bit budget. This implies that SNNs may have an advantage over ANNs in scenarios of ultra-low-energy computation, which is in line with previous studies [deng2020rethinking, he2020comparing].
@@ 6.1 Maintaining the neuromorphic computing nature
Exploiting the spike perceptrons. Temporal squeezing itself is consistent with neurodynamics as it is essentially composed of a signal accumulation [gerstner2014neuronal] and a division operation. A similar accumulation mechanism can also be found in the current accumulation stage of spiking neuron, such as LIF and IF. Particularly, SpikeConverter neuron [liu2022spikeconverter] separates this accumulation from the spiking neuron modeling and uses a special computing pipeline to handle this accumulation. Their accumulation is identical to ours, which is proven feasible in asynchronous computing. In addition, such accumulation has also been realized and implemented by neuromorphic circuits to process the DVS signals and generate the dynamic datasets [amir2017low, lin2021imagenet].
@@
Using the two-stage pipeline. Although we have theoretically elaborated on the harmlessness of temporal squeezing to the asynchronous nature in the above, one may argue that the entire network’s spike flow could be interrupted by the squeezing stage, leading to longer inference latency. This concern has already been discussed and resolved by [liu2022spikeconverter], where the two-stage pipeline is adopted. The first stage is spike accumulation, i.e., the temporal squeezing, the second stage is spike emission. These two separate stages can overlapped inter-layer. Therefore, using temporal squeezing only introduces very limited latency increase.
6.2 Hardware considered energy merit analysis
[davidson2021comparison] explains that SNNs are not inherently more energy-efficient than ANNs in all scenarios, detailing a specific condition where SNN energy becomes lower. This specific condition demands "the expected number of spikes per output" to be less than 1.72 on their hardware implementation.
As shown in Table 1 and Table 2, our the temporal lengths of our SNNs are mostly learned to 1. Clearly, "the expected number of spikes per output" is less than 1.72. However, we need to consider a more rigorous situation where "the expected number of spikes per output" is actually "the expected number of non-zero bits in activation states per output", because our spiking neuron generates a multi-bit spike at a time-step while [davidson2021comparison] generates a 1-bit spike at a time-step. Thus, we can have:
activation states $=T\cdot S,$ (27)
$\begin{split}\text{Exp(act.)}&=\text{Exp}(T\cdot S\cdot fr_{s})\\ &\approx T\cdot S\cdot\overline{fr}_{s}.\end{split}$ (28)
Here, Exp(act.) abbreviates "the expected number of non-zero bits in activation states per output". Given Equation 8 and Equation 9, we can estimate Exp(act.) via:
$\begin{split}\text{Exp(act.)}&\approx T\cdot S\cdot\overline{fr}_{s}\\ &=T\cdot S\cdot\frac{\text{NS-ACE}}{\text{S-ACE}}.\end{split}$ (29)
Given Table 2, we can get our SNNs’ Exp(act.)s are 0.55, 1.23, and 1.22, respectively. We also record their "average numbers of non-zero bits in activation states per output" through sampling the data from a random mini-batch of ImageNet. The recorded "average numbers of non-zero bits in activation states per output" are 0.57, 1.28, and 1.22, respectively.
Both of the estimated and recorded results are all less than 1.72, which implies our SNNs should be more energy-efficient than ANNs on the hardware implementation of [davidson2021comparison].
6.3 Reusable neural integration unit design
Figure 5: Bit-serial style neural integration unit that can handle various W/S/T combinations. Psum represents the input current. "Reg." denotes register. "Div." is divider. For simplicity, we set S=2 and T=2.
One may argue that [shen2024conventional] proved the "Bit Budget" concept based on the static/unconfigurable neural integration unit. Our proposed layer-wise bit allocation would incur massive hardware resource waste if the neural integration units were dedicated to specific W/S/T combinations and could not be reused across different layers. Thus, the "Bit Budget" concept may become ineffective.
Here, we draw the design philosophy of bit-serial hardware [sharma2018bit, albericio2017bit, sharify2019laconic, sharify2018loom] to provide a highly reusable neural integration unit that can handle any W/S/T combination, as shown in Figure 5. For simplicity and flexibility, we utilize two counters to implement the two finite state machines (FSMs) for the timestep loop (T) and the spike bit loop (S), respectively. Thus, the FSMs become easy to configure, without any configuration overhead and resource waste. Simply adding some necessary peripheral logic circuits to generate the enabling signals will make the local controller functional. For different T and S combinations, we only need to reset the counters’ count limits. Another design core is the bit-serial style divider [parhami1999computer] that outputs one bit value at one cycle. Based on this, the whole unit only handles the computation of one bit value at one cycle. In short, one bit in, and one bit out. As for the following matrix multiplications between spikes and weights, e.g., convolutions, employing one of those advanced mixed-precision architectures [tahmasebi2024flexibit, luo2023deepburning, li2022bit, zhang2024mixpe] would effectively and efficiently suffice.
Therefore, with our bit-serial style design, any W/S/T combination can be processed by the same neural integration unit. The computations of different layers can be mapped to the same processing-element (PE) array. Thus, the potential hardware utilization issue can be resolved.
Figure 6: Relationship between bit budgets and the energy consumption of individual synaptic operation.
6.4 Energy consumption analysis of individual synaptic operation
Based on our bit-serial style hardware implementation, we can also prove the same linear relationship between bit budget and energy consumption for individual synaptic operations on the FPGA platform, as [shen2024conventional] proved before.
As shown in Figure 6, we conduct the analogous energy analysis as [shen2024conventional]. The linear relationship between bit budgets and synaptic energy is very intuitive. Therefore, we can conclude that the "Bit Budget" concept should be consistently effective for our layer-wise bit-adaptive SNNs.